OpenStudy (anonymous):

Solve 11.32x – 1 = 15.7 x ≈ 1.12 x ≈ 0.57 x ≈ 0.94 x ≈ 1.07

OpenStudy (callisto):

Add 1 to both sides. 11.32x – 1 = 15.7 11.32x – 1 +1 = 15.7 +1 11.32x = 16.7 Divide both sides by 11.32. What do you get?

OpenStudy (anonymous):

add 1 to both sides to cancel 1 then divide both sides by 11.32 to isolate x 11.32x - 1 +1 = 15.7 +1 11.32x = 16.7 11.32/11.32x = 16.7/11.32 x?

OpenStudy (anonymous):

alternate forms 11.32 x-1. = 15.7

OpenStudy (anonymous):

x~~1.47527

OpenStudy (anonymous):

i think ur question is wrong

OpenStudy (callisto):

@best.shakir Don't give direct answer without any steps provided and interaction with the asker. Thanks.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The answer is wrong possible answers were x = 1.12 x = .57 x = .94 x = 1.07